OpenStudy (anonymous):

-8=1/3x+x show the steps

OpenStudy (anonymous):

to be honest it's confusing... if you mean that -8=(1/3)*x+x then -8=1/3*x+x=4/3*x x=-6 or if you mean -8=1/(3*x)+x then -8x=1/3+x^2 3x^2+24x+1=0 then x=[(-12\pm \sqrt{141})/3\]

jhonyy9 (jhonyy9):

\[-8=1/3x +x ---- -24x=1+3x ^{2} ----- 3x ^{2}+24x+1=0\]

jhonyy9 (jhonyy9):

\[3x ^{2} +24x +1=0\]